What is an equation for surface area of a box with 12 inches long, 12 inches wide and 12 inches tall?
Found 2 solutions by Alan3354, Parasity:
Answer by Alan3354(69443)

What is an equation for surface area of a box with 12 inches long, 12 inches wide and 12 inches tall?
----------------
Each face is 12 by 12, there are 6 faces.
Area = 6*12*12 sq inches = 864 sq inches
---------------------------
That's the solution. The formula for sufrace area of a "box" is
2*(l*w + l*h + w*h)

Answer by Parasity(1)

To calculate the surface are of a cube, find the surface area of one side and multiply by 6. The surface area of any side is the length of a side squared.

So: SA = 2(length*width) + 2(height*width) + 2(length*height)
2(12*12)+2(12*12)+2(12*12)
2(144)+2(144)+2(144)
288+288+288 = 864 inches.
So the answer is: 864 inches.